8. A probability density function (PDF) is given by: f(x)-k(8x-x2) for 0cx<8 What value of k will make this a PDF? 9. A probability density function (PDF) is given by: f(x)-e.( 4) for x>a What value of a will make this a PDF? 10. A probability density function (PDF) is given by: f(x)-1.5x2 for -acx<a What value of a will make this a PDF?
Answer #1

8) Given the PDF fleft ( x ight )=kleft ( 8x-x^2 ight );0<x<8 . The conditions for a function to be PDF are

fleft ( x ight )geqslant 0 extup{ and } int_{0}^{8}fleft ( x ight )dx=1. Here

int_{0}^{8}fleft ( x ight )dx=int_{0}^{8}kleft ( 8x-x^2 ight )dx int_{0}^{8}fleft ( x ight )dx=kleft ( 8x-x^2 ight ]_{0}^{8} int_{0}^{8}fleft ( x ight )dx=kleft ( 256-512/3 ight ]

Now setting conditions,

kleft ( 256-512/3 ight ]=1Rightarrow{color{Blue} k=3/256}

9) Given the PDF fleft ( x ight )=e^{x-4};x>a . The conditions for a function to be PDF are

f (x) 0 and f (x) dr1. Here

f(x) dx= | e-(-4) dx 1

10) Given the PDF f(x) = 1.5广-a < x < a . The conditions for a function to be PDF are

f (x) 0 and f (x) dr1. Here

f(x) dx=| 1.512 dr -a -a 1.5x2dr= [0.5레-a -0 Cl